# Local consistency and axioms of functional dependence

### v1

@@ 3.1 $K$-realisability of Contextual Families
Item 1 is immediate, item 2 follows by local consistency, and item 3 follows by the construction of $M$. Since $\leq$ is transitive and reflexive, we obtain #$\#(M_{i})\leq\#(M_{i+1})$. Also, for the first inequality in (2), we obtain $M_{0}\cup\{s\}\subseteq\mathcal{R}^{\prime}_{0,+}$, #$\#(\mathcal{R}^{\prime}_{0,+})=\#(\mathcal{R}^{\prime}_{1,-})$, $\mathcal{R}^{\prime}_{1,-}=M_{1}$, where $\mathcal{R}^{\prime}_{0,+},\mathcal{R}^{\prime}_{1,-}$ are defined otherwise as $\mathcal{R}_{0,+},\mathcal{R}_{1,-}$, except that $M_{0}$ is replaced with $M_{0}\cup\{s\}$. This concludes the proof of the $\Rightarrow$” direction
( $\Leftarrow$ ) Suppose $\mathrm{OPGraph}(\mathcal{R})$ has an edge cycle cover. Then, each assignment $s\in\bigcup{\mathrm{Supp}(\mathcal{R})}$ belongs to a cycle in $\mathrm{OPGraph}(\mathcal{R})$. In particular, there must be a simply cyclic contextual family $\mathcal{T}_{s}$ such that $s\in\bigcup{\mathrm{Supp}(\mathcal{T}_{s})}\subseteq\bigcup{\mathrm{Supp}(\mathcal{R})}$. Then, Equation 1 entails $w\mathcal{T}_{s}$ is a contextual $K$-family, for any $w\neq 0$ (which exists since $K$ is non-trivial). Since contextual $K$-families over $\mathcal{C}$ are closed under addition (Prop. 3), the sum $\sum_{s\in\bigcup{\mathrm{Supp}(\mathcal{R})}}w\mathcal{T}_{s}$ is a contextual $K$-family. Its support is $\mathcal{R}$, for $K$ is positive. Hence the $\Leftarrow$” follows. ∎
Consider our running example in light of this theorem:
Example 9.
[…]
3.2 Contextual $\mathbb{R}_{\geq 0}$-realisability
Let us next briefly consider the special case of $\mathbb{R}_{\geq 0}$-realisability. As a consequence of the previous section, we observe that a contextual $\mathbb{R}_{\geq 0}$-family decomposes to its cycles. Moreover, we establish that a contextual family is the support of some contextual $\mathbb{R}_{\geq 0}$-family if and only if it is the support of some contextual $\mathbb{N}$-family.
Corollary 10.
Each contextual $\mathbb{R}_{\geq 0}$-family over a chordless-cycle context set is a non-negative combination of its cycles.
@@ 3.2 Contextual $\mathbb{R}_{\geq 0}$-realisability
Let $\mathcal{R}$ be a contextual $\mathbb{R}_{\geq 0}$-family, with a multiplicity mapping #$\#\colon\bigcup{\mathrm{Supp}(\mathcal{R})}\to\mathbb{R}_{>0}$. Assume $\mathcal{R}$ is non-empty; otherwise the claim holds trivially. Theorem 8 entails that there is a non-empty set of assignments $S\subseteq\bigcup{\mathrm{Supp}(\mathcal{R})}$ forming a simple cycle in $\mathrm{OPGraph}(\mathcal{R})$. Let $w=\min\{\#(s)\mid s\in S\}$. Define the updated family $\mathcal{R}^{\prime}\coloneqq\mathcal{R}-wS$, where subtraction is defined in the obvious way. Then, $\mathcal{R}^{\prime}$ is a contextual $\mathbb{R}_{\geq 0}$-family. We repeat this process recursively, setting $\mathcal{R}\coloneqq\mathcal{R}-wS$ at each step, until the family becomes empty. Since $S$ is removed entirely at each step (i.e., it no longer appears in the support), no simple cycle $S$ is used more than once, meaning that the process will terminate. This leads to a decomposition of the form $\mathcal{R}=w_{1}S_{1}+\dots+w_{\ell}S_{\ell}$. ∎
This concludes our discussion of context sets that form chordless cycles. A natural follow-up question is whether a similar graph-theoretic characterisation exists for $\mathbb{R}_{\geq 0}$-realisability over arbitrary context sets. The next example suggests that the answer is negative. We now write $x_{1}\dots x_{n}\mapsto a_{1}\dots a_{n}$ for the assignment that maps $x_{1},\dots,x_{n}$ respectively to $a_{1},\dots,a_{n}$.
Example 11.
Consider the contextual family represented in Figure 3. We note that restricted to $\{ab,bc,ca\}$ the corresponding overlap projection graph has an edge cover, and thus the restriction is $\mathbb{R}_{\geq 0}$-realisable by Theorem 8. Similar reasoning applies to the restriction on $\{ab^{\prime},b^{\prime}c,ca\}$. However, over the full context set, $\mathbb{R}_{\geq 0}$-realisability does not hold anymore. Assume toward a contradiction that this is not the case. Let $S$ be the set of assignments represented in Figure 3. Then, Proposition 12 yields a mapping $\pi\colon S\to\mathbb{N}$ which forms a contextual $\mathbb{N}$-relation and is such that $\pi(bc\mapsto 01)=m\geq 1$. By local consistency, $\pi(ca\mapsto 10)=n\geq m$, and furthermore $\pi(ca\mapsto 10)=\pi(ab^{\prime}\mapsto 00)=\pi(b^{\prime}c\mapsto 00)=\pi(ca\mapsto 01)=\pi(ab\mapsto 11)=\pi(bc\mapsto 11)$. Thus
$\pi(bc\mapsto 01)+\pi(bc\mapsto 11)=n+m\neq n=\pi(ca\mapsto 10),$
violating local consistency.
$\begin{array}[t]{cccccc}\begin{array}[t]{cc}a&b\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}b&c\\ \hline\cr\hline\cr 0&0\\ 0&1\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}c&a\\ \hline\cr\hline\cr 0&1\\ 1&0\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}a&b^{\prime}\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}b^{\prime}&c\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}\end{array}$
Figure 3: Non-realisable contextual family.
Proposition 12.

### v2

@@ Appendix A Full proofs
Item 1 is immediate, item 2 follows by local consistency, and item 3 follows by the construction of $M$. Since $\leq$ is transitive and reflexive, we obtain $(M_{i})\leq\#(M_{i+1})$. Also, for the first inequality in (2), we obtain $M_{0}\cup\{s\}\subseteq\mathcal{R}^{\prime}_{0,+}$, $(\mathcal{R}^{\prime}_{0,+})=\#(\mathcal{R}^{\prime}_{1,-})$, $\mathcal{R}^{\prime}_{1,-}=M_{1}$, where $\mathcal{R}^{\prime}_{0,+},\mathcal{R}^{\prime}_{1,-}$ are defined otherwise as $\mathcal{R}_{0,+},\mathcal{R}_{1,-}$, except that $M_{0}$ is replaced with $M_{0}\cup\{s\}$. This concludes the proof of the $\Rightarrow$” direction
( $\Leftarrow$ ) Suppose $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$ is edge-cyclic. Then, each assignment $s\in\bigcup{\mathrm{Supp}(\mathcal{R})}$ belongs to a cycle in $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$. In particular, there must be a cyclic contextual family $\mathcal{T}_{s}$ such that $s\in\bigcup{\mathrm{Supp}(\mathcal{T}_{s})}\subseteq\bigcup{\mathrm{Supp}(\mathcal{R})}$. Then, Lemma 27 entails $w\mathcal{T}_{s}$ is a contextual $K$-family, for any $w\neq 0$ (which exists since $K$ is non-trivial). Since contextual $K$-families over $\mathcal{C}$ are closed under addition (Prop. 4), the sum $\sum_{s\in\bigcup{\mathrm{Supp}(\mathcal{R})}}w\mathcal{T}_{s}$ is a contextual $K$-family. Its support is $\mathcal{R}$, for $K$ is positive. Hence the $\Leftarrow$” follows. ∎
Full proof for Proposition 23.
Proposition 23.
The problem of deciding whether $x_{1}\rightarrow x_{n}$ can be derived from a set of CDs and unary FDs $\Sigma$ by a single application of the contextual chain rule is $\mathrm{NL}$-complete.
@@
To verify that the problem is in $\mathrm{NL}$, we need to show that it can be decided in nondeterministic logarithmic space. To that end, it is crucial that each of the five conditions of the contextual chain rule can be checked step by step, without the need of keeping the previous steps in memory. Given a set of CDs and unary FDs $\Sigma$ and two variables $x_{1}$ and $x_{n}$, the algorithm to solve the aforementioned decision problem works as follows:
First, we nondeterministically guess $c_{1}$ and check whether $c_{1}\rightarrow x_{n}\in\Sigma$. Afterwards, for all $i\in(2,\dots,\mbox{n-1})$, we guess $c_{i}$ and $x_{i}$ and check if $x_{i-1}\rightarrow x_{i},c_{i}\rightarrow x_{n},x_{i-1}c_{i-1}x_{i},c_{i-1}x_{i}c_{i},c_{i-1}c_{i},x_{n}\in\Sigma$. Finally, we check whether $x_{n-1}c_{n-1}x_{n}\in\Sigma$.
If all those checks are passed, we know that there exist $x_{2},\dots,x_{-1}$ and $c_{1},\dots,c_{n-1}$ such that the conditions of the rule are met and thus $x_{1}\rightarrow x_{n}$ can be derived from $\Sigma$ by a single application of the contextual chain rule. Furthermore, we only ever had to keep four non-input values, i.e., $x_{i-1},x_{i},c_{i-1}$ and $c_{i}$, in memory, for which logarithmic space is sufficient.
To show $\mathrm{NL}$-hardness, we reduce from the well-known $\mathrm{NL}$-complete graph reachability problem (?): for a graph $G$ and two nodes $s,t$, $(G,s,t)\in\mathrm{REACH}$ if and only if there is a path from $s$ to $t$ in $G$.
Note that the first condition of the contextual chain rule essentially encapsulates reachability from $x_{1}$ to $x_{n}$ in the dependency graph of $\Sigma$. The four further conditions only add the requirement that all those “edges” need to be “witnessed” by a variable $c_{i}$. If we transform a graph into a set of dependence atoms where each “edge” is trivially “witnessed”, we end up with a plain reachability problem again.
Given $G=(V,E)$ and two nodes $s$ and $t$, we produce the unary FDs $\{u\rightarrow v\mid(u,v)\in E\}\cup\{c\to t\}$ and the CDs $\{uvc\mid(u,v)\in E\}$. In the resulting set $\Sigma$, there is a chain of FDs from $s$ to $t$, if and only if there is a path from $s$ to $t$ in $G$. And since all edges are witnessed, the four conditions beyond the first are trivially met. Finally, this function is a logspace reduction, since only one edge at a time needs to be kept in memory. In total, this yields $\mathrm{NL}$-completeness of one application of the chain rule. ∎
Full proof for Theorem 24.
Theorem 24.
The problem whether a unary FD $y_{1}\rightarrow y_{2}$ can be derived from a set $\Sigma$ of FDs and CDs by using the reflexivity rule, the cycle rule and the contextual chain rule is decidable in polynomial time.